# TRecViT: A Recurrent Video Transformer

### v1

@@ 3.2 Gated LRUs for Video
We adopt the gated variant of the LRU [9] to design our proposed block for video modelling.
Let $X\in[0,1]^{T\times H\times W\times 3}$ be an RGB video with $T$ frames and $H\times W$ pixels. The video frames are split into $N$ non-overlapping patches $p_{t}^{k}$ of size $n\times n\times 3$, with $t\in\{1,T\}$ and $k\in\{1,N\}$. Let $x_{t}^{k}$ be the tokens obtained after the linear projection of the patches and the addition of the spatial positional encoding, with token size $1\times 1\times d$, where $d$ is the token feature dimension. Each LRU operates over a temporal tube $\{x_{t}^{k}|t=\overline{1,T}\}$, following the equations below (we drop the $k$ spatial index for clarity):
$i_{t}$ $=$ $\sigma(W_{x}x_{t}+b_{x}),\quad\,\,\,{\color[rgb]{.5,.5,.5}\text{\emph{ input gate}}}$ (1)
@@
$r_{t}$ $=$ $\sigma\left(W_{\lambda}x_{t}+b_{\lambda}\right),\quad{\color[rgb]{.5,.5,.5}\text{ \emph{recurrence gate}}}$ (2)
$\lambda_{t}$ $=$ $\sigma(\lambda)^{\text{C}\cdot r_{t}},$ (3)
$h_{t}$ $=$ $\lambda_{t}\odot h_{t-1}+\sqrt{1-\lambda_{t}^{2}}\odot(i_{t}\odot x_{t}).$ (4)
@@
where $h_{t}\in\mathbb{R}^{d}$ is the state of the LRU, $\lambda_{t}\in\mathbb{R}^{d}$ is a vector containing the eigenvalues of the (diagonal) recurrence matrix [^2] Similar to [9], we implement the recurrence weights $\lambda_{t}$ as ${\exp(-C\cdot\text{softplus}(\lambda)\cdot r_{t})}$, which is mathematically equivalent but numerically more stable., $i_{t}\in\mathbb{R}^{d}$ is the input gate controlling whether $x_{t}\in\mathbb{R}^{d}$ is integrated within the state $h_{t}$ of the LRU or not, and $r_{t}\in\mathbb{R}^{d}$ is the recurrence gate. The weights and biases of the LRU ($W_{x}\in\mathbb{R}^{d\times d}$, $W_{\lambda}\in\mathbb{R}^{d\times d}$, $b_{x}\in\mathbb{R}^{d}$, $b_{\lambda}\in\mathbb{R}^{d}$) are initialized using LeCun init [26].
@@
The (learnable) recurrence weights $\lambda$ are passed through a sigmoid function to ensure they are between 00 and $1$, and are initialised such that $\sigma(\lambda)$ is sampled uniformly in $[\lambda_{\min},\lambda_{\max}]$. These recurrent weights are raised to the power $\text{C}\cdot r_{t}$, which effectively acts as a gate controlled by $r_{t}$ given in equation (2). $r_{t}$ is defined as a linear projection, with parameters $W_{\lambda}$ and $b_{\lambda}$, followed by a sigmoid function to ensure again the range $[0,1]$. By raising element-wise $\sigma(\lambda)$ to $r_{t}$, the effective recurrence weight at some position $j$ can change between the $j$-th entry of $\sigma(\lambda)$ when the corresponding gate entry is $1$ and $1$ when the gate entry is 0.
@@
The additional constant coefficient $\text{C}\in\mathbb{R}$, typically set to 8 as in [9], increases the range to be between $\sigma(\lambda)^{\text{C}}$ to $1$, providing additional flexibility. E.g. if $\sigma(\lambda)$ is $0.9$ and we set $\text{C}=8$, we extend the range from $[0.9,1]$ to $[0.43,1]$. More importantly, we change the learning dynamics (e.g. gradient norms) and resolution we have over the range during learning. Specifically, for $x_{t}$ in some fixed interval and similar magnitude $W_{\lambda}$, as it is the case at initialisation, a higher value of C implies $\lambda_{t}$ will concentrate more towards the edges of the range. Note also that this is the dynamic range in which the recurrent weights can vary during inference as a function of the input tokens.
@@
In [9], the authors found that setting $\lambda_{\min}=0.9$ and $\lambda_{\max}=0.999$ leads to the best results. An eigenvalue of $0.9$ implies that it will take at least $10$ time steps for the information to decay to roughly $35\%$ of its magnitude, while for an eigenvalue of $0.999$ it will take $1000$ time steps to decay by the same amount. When using the same range for video modelling, we observed that the eigenvalues are pushed significantly towards $\lambda_{\min}$ during training, with a small number of eigenvalues becoming smaller than $\lambda_{\min}$; see Figure 2. We experimented with extending the range and obtained better results with $\lambda_{\min}=0.6$. This leads to faster decay of information initially and might reflect the importance for videos of having enough recurrent units focused on short term information, in order to disentangle fast changing dynamics from slow ones.
@@
Finally, note that when diagonalising the recurrence matrix, the eigenvalues $\lambda$ could, in theory, have complex values. We conducted experiments using complex eigenvalues, but we did not see improvements compared to using only real eigenvalues. The same observation was made in  [9, 16] as well.
(a) Memory comparison
(b) FLOPs comparison
Figure 3: Our model demonstrates increasingly greater memory and compute savings compared to ViViT baselines as the number of frames increases. For clarity, TRecViT’s peak memory (left figure) goes from about 4G for 8 frames to 22.4G for 64 frames, but this increase is dwarfed by ViViT’s increase, hence TRecViT line appears almost horizontal
@@ 3.3 Video block based on gated LRU
We use the gated LRU in a similar block structure as the one employed in [9], see Figure 1b. Given a 1D input (temporal tube), the block first applies a normalisation layer, then the signal is routed on two different paths. On the first one, it gets linearly projected to same dimensionality $d$ and then the GeLU activation is applied. On the other path, the signal is also linearly projected to the same dimensionality $d$, then we apply a 1D convolution followed by the gated LRU described in equation (4). The output of the LRU and the GeLU branch are element-wise multiplied and then linearly projected to the same dimension $d$. Note that, in line with [9], we use a separable convolution, which allows mixing information only over time, not over channels. We sweep the width of the convolutional kernel and find that a window of 2 is enough compared to [9] which used 4. Also, different from  [9], we do not use an MLP block after the LRU for feature mixing. We apply the MLP after the self-attention block, as done in ViT.
Given the diagonal form of the recurrence, on device, the gated LRU computations are memory-bound, i.e. the data transfer takes longer than the actual computations done on that data. Similar to [9] we use a specialised Pallas [6] kernel that minimizes the number of bytes that need to be moved between HBM and VMEM (the Vector Processing Unit’s cache). The parameters added by the linear projections within the block, as well as the parameters of the convolution and the LRU, are learned.
@@ 4.1 Self-supervised pre-training
We follow the same recipe as in the original VideoMAE paper [40]. Specifically, we use tube masking where a 2D random mask is generated and repeated for all the frames in the video. For our architecture, this is equivalent to dropping temporal LRUs. The training objective is simply L2 reconstruction error of the entire frames. We sweep the value of the masking ratio and we find that 0.90 leads to best performance on downstream tasks. When using the pre-trained representations for downstream tasks, we keep all the tokens of the video and we add a decoder or readout head that is fine-tuned for the respective tasks.
4.2 Memory footprint and FLOPs
We compare the memory footprint and the number of FLOPs of TRecViT against ViViT baselines, see Figure 3. The profiling results are obtained by cost and memory analysis of lowered Jax HLO on CPU backend to be aligned with the theoretical numbers [2]. We consider as input a video of size $224\times 224$ and we vary the length of the video to analyse the savings provided by our architecture as the length of the video increases. Although in number of parameters for TRecViT is in between ViViT-B and ViViT-L (90M $>$ 109M $>$ 320M), the peak memory and number of flops for TRecViT are significantly lower as the number of frames increases, e.g. at 32 frames (the number of frames typically used in video classification experiments), TRecViT’s peak memory is $\sim$12$\times$ smaller than that of ViViT-L and the FLOPs count is 5×5$\times$5 × lower. When going to 64 frames, the peak memory is $\sim$24$\times$ smaller and FLOPs count is 8×8$\times$8 × lower.
@@ 5 Experiments
We present results for supervised video classification and self-supervised masked auto-encoding with frozen representations evaluated on two downstream tasks: video classification and point tracking. To analyse the memory capabilities of our model, we also include a reconstruction task of frames seen in the distant past. Using the same task, we study the generalisation capabilities to longer sequences than seen during training. We follow the ViT scaling configurations and, unless otherwise stated, we use the Base version for our model for all our experiments. We specify the number of parameters for all models considered in our experiments, and we include in the supplementary material all the training hyperparameters and data augmentations used in all experiments.
5.1 Supervised video classification
@@
Datasets: We use large-scale real-world datasets for the supervised video classification task. Kinetics400 [7] contains 241,512 videos [^3] Kinetics is a dynamic dataset (videos may be removed from YouTube). Our current version has 241,512 videos, compared to 267,000 videos reported in [1], so a decrease of almost 10%, noticeable in the final performance. across train, validation, and test splits, 10s-long (25fps), spanning 400 classes. This dataset is known to require modelling appearance for successful action recognition. To challenge our model’s capability of understanding motion, we also use SSv2 dataset [15], which contains 220,847 shorter videos (2-6s long), sampled at 12fps, representing 174 classes. This dataset includes actions that differ in finer motion-related details, requiring a deeper temporal understanding, e.g. pouring something into something vs pretending to pour something into something.
Baselines: We use ViViT [1] as our main baseline. We consider the full self-attention version, which patchifies and flattens the entire video, prepends a video class token, then runs self-attention blocks. We also consider the factorised encoder version (ViViT FE), which runs a ViT image model over all the frames, and uses temporal self-attention blocks to integrate the information over time. Finally, we also consider a baseline that uses only LRU recurrent and MLP blocks, configured similar to VideoMamba [27], i.e. it does not use self-attention blocks, denoted PureLRU. Similar to ViViT, this model first patchifies and flattens the video, prepends a class token, then applies a sequence of recurrent blocks. All baselines use learnt spatio-temporal positional encoding, whereas the proposed TRecViT uses only spatial positional encoding as the temporal dimension is implicitly modelled through its recurrence.
@@
Results: We include results for training from scratch or using Imagenet pre-trained weights to initialise the weights of the ViT blocks. Figure 4 shows a first comparison between TRecViT and the above baselines, with all models being trained from scratch on supervised classification on SSv2. We consider the Small version for all models as the larger Base version shows stability issues when trained from scratch, as reported in other works as well [27, 1]. As expected, the performance on this challenging dataset when training from scratch is far from SOTA, but it clearly shows that the proposed factorisation has superior video modelling capabilities compared to baselines, ViViT-S with full self-attention being the closest competitor. PureLRU’s performance is very poor, which is in line with the findings of other works (e.g. VideoMamba) who report that bidirectional (non-causal) processing of the input is needed for good performance.
We report further results comparing against ViViT-B and ViViT-L with full self-attention when using Imagenet pre-trained weights; see Table 1 for SSv2 results and Table 2 for Kinetics400 results. We can observe that our model achieves better performance compared to ViViT baselines on SSv2, but it is slightly below ViViT-L on Kinetics400. This result could reflect the difference between the two datasets mentioned above: outperforming ViViT-L on SSv2 suggests that TRecViT is superior at modelling motion compared to ViViT, but on Kinetics where the appearance is enough for successful classification, both models are on par. We consider this to be a strong positive result for our model given that it has about 3x less parameters compared to ViViT-L and significantly lower FLOPs count and memory footprint as shown in Figure 3.

### v2

@@ 3.1 Gated LRUs for Video
We adopt the gated variant of the LRU (De et al., 2024) to design our proposed block for video modelling. Although LRUs have been proposed for language, we hypothesise that they can effectively model video signals, given the continuous nature of the video and LRUs’ inspiration from time continuous systems. We run extensive analysis and ablations for the different components and hyperparameters of the gated LRU block to find a configuration that works well for video; see analysis below and ablations in section 5.1.
The gated LRU operation is described by the below equations:
$i_{t}$ $=$ $\sigma(W_{x}x_{t}+b_{x})\quad\,\,\,{\color[rgb]{.5,.5,.5}\text{\emph{ input gate}}}$ (1)
@@
$r_{t}$ $=$ $\sigma\left(W_{\lambda}x_{t}+b_{\lambda}\right)\quad{\color[rgb]{.5,.5,.5}\text{ \emph{recurrence gate}}}$ (2)
$\lambda_{t}$ $=$ $\sigma(\lambda)^{\text{C}\cdot r_{t}}$ (3)
$h_{t}$ $=$ $\lambda_{t}\odot h_{t-1}+\sqrt{1-\lambda_{t}^{2}}\odot(i_{t}\odot x_{t})$ (4)
@@
where $h_{t}\in\mathbb{R}^{d}$ is the state of the LRU, $\lambda_{t}\in\mathbb{R}^{d}$ is a vector containing the eigenvalues of the (diagonal) recurrence matrix [^1] Similar to (De et al., 2024), we implement the recurrence weights $\lambda_{t}$ as $\exp(-C\cdot\text{softplus}(\lambda)\cdot r_{t})$}, which is mathematically equivalent but numerically more stable., $i_{t}\in\mathbb{R}^{d}$ is the input gate controlling whether $x_{t}\in\mathbb{R}^{d}$ is integrated within the state $h_{t}$ of the LRU or not, and $r_{t}\in\mathbb{R}^{d}$ is the recurrence gate. The weights and biases of the LRU ($W_{x}\in\mathbb{R}^{d\times d}$, $W_{\lambda}\in\mathbb{R}^{d\times d}$, $b_{x}\in\mathbb{R}^{d}$, $b_{\lambda}\in\mathbb{R}^{d}$) are initialized using LeCun init (LeCun et al., 2012).
@@
The (learnable) recurrence weights $\lambda$ are passed through a sigmoid function to ensure they are between 0 and $1$, and are initialised such that $\sigma(\lambda)$ is sampled uniformly in $\lambda_{\min},\lambda_{\max}]$. These recurrent weights are raised to the power $\text{C}\cdot r_{t}$, which effectively acts as a gate controlled by $r_{t}$ ( equation equation 2). $r_{t}$ is defined as a linear projection, with parameters $W_{\lambda}$ and $b_{\lambda}$, followed by a sigmoid function to ensure again the range $[0,1]$. By raising element-wise $\sigma(\lambda)$ to $r_{t}$, the effective recurrence weight at some position $j$ can change between the $j$-th entry of $\sigma(\lambda)$ when the corresponding gate entry is $1$ and $1$ when the gate entry is 0.
@@
The additional constant coefficient $\text{C}\in\mathbb{R}$, typically set to 8 as in (De et al., 2024), increases the range to be between $\sigma(\lambda)^{\text{C}}$ to $1$, providing additional flexibility. E.g. if $\sigma(\lambda)$ is $0.9$ and we set $\text{C}=8$, we extend the range from $[0.9,1]$ to $[0.43,1]$. More importantly, we change the learning dynamics (e.g. gradient norms) and resolution we have over the range during learning. Specifically, for $x_{t}$ in some fixed interval and similar magnitude $W_{\lambda}$, as it is the case at initialisation, a higher value of C implies $\lambda_{t}$ will concentrate more towards the edges of the range. Note also that this is the dynamic range in which the recurrent weights can vary during inference as a function of the input tokens.
@@
In (De et al., 2024), the authors found that setting $\lambda_{\min}=0.9$ and $\lambda_{\max}=0.999$ leads to the best results. An eigenvalue of $0.9$ implies that it will take at least $10$ time steps for the information to decay to roughly $35\%$ of its magnitude, while for an eigenvalue of $0.999$ it will take $1000$ time steps to decay by the same amount. When using the same range for video modelling, we observed that the eigenvalues are pushed significantly towards $\lambda_{\min}$ during training, with a small number of eigenvalues becoming smaller than $\lambda_{\min}$; see Figure 2. We experimented with extending the range and obtained better results with $\lambda_{\min}=0.6$; see Table 1. This leads to faster decay of information initially and might reflect the importance for videos of having enough recurrent units focused on short term information, in order to disentangle fast changing dynamics from slow ones.
@@
Finally, note that when diagonalising the recurrence matrix, the eigenvalues $\lambda$ could, in theory, have complex values. We conducted experiments using complex eigenvalues, but we did not see improvements compared to using only real eigenvalues. The same observation was made in  (De et al., 2024; Gu and Dao, 2023) as well.
@@ 3.2 Video block based on gated LRU
We use the gated LRU in a similar block structure as the one employed in (De et al., 2024), see Figure 1b. Given a 1D input (temporal tube), the block first applies a normalisation layer, then the signal is routed on two different paths. On the first one, it gets linearly projected to same dimensionality $d$ and then the GeLU activation is applied. On the other path, the signal is also linearly projected to the same dimensionality $d$, then we apply a 1D convolution followed by the gated LRU described in equation 4. The output of the LRU and the GeLU branch are element-wise multiplied and then linearly projected to the same dimension $d$. Note that, in line with (De et al., 2024), we use a separable convolution, which allows mixing information only over time, not over channels. We sweep the width of the convolutional kernel and find that a window of 4 gives the best results, similar to (De et al., 2024). Different from  (De et al., 2024), we do not use an MLP block after the LRU for feature mixing. We apply the MLP after the self-attention block, as done in ViT.
Given the diagonal form of the recurrence, on device, the gated LRU computations are memory-bound, i.e. the data transfer takes longer than the actual computations done on that data. Similar to (De et al., 2024) we use a specialised Pallas (Bradbury et al., 2018) kernel that minimizes the number of bytes that need to be moved between HBM and VMEM (the Vector Processing Unit’s cache). The parameters added by the linear projections within the block, as well as the parameters of the convolution and the LRU, are learned.
@@ 4.1 Self-supervised pre-training
We follow the same recipe as in the original VideoMAE paper (Tong et al., 2022). Specifically, we use tube masking where a 2D random mask is generated and repeated for all the frames in the video. For our architecture, this is equivalent to dropping temporal LRUs. The training objective is simply L2 reconstruction error of the entire frames. We sweep the value of the masking ratio and we find that 0.90 leads to best performance on downstream tasks. When using the pre-trained representations for downstream tasks, we keep all the tokens of the video and we add a decoder or readout head that is fine-tuned for the respective tasks.
4.2 Computational complexity and efficiency analysis
Figure 3: Left: Memory comparison; Right: FLOPs comparison. Our model demonstrates increasingly greater memory and compute savings compared to ViViT full self-attention as the number of frames increases. The efficiency of ViViT factorised self-attention is very similar to TRecViT’s, but its performance is significantly lower. The factorised encoder variant of ViViT-L is very efficient, but it performs well mainly on sparse tasks.
We discuss here the computational complexity of the proposed architecture and its efficiency compared to different variants of ViViT. Our model’s parameter count ( [^1] M) falls between that of ViViT-B (90M) and ViViT-L (310M). We focus on ViViT-L as the main point of comparison in our efficiency and SOTA analysis (Section 5.2) as this is the strongest representative of the ViViT family reported in the original ViViT paper. Furthermore, since our architecture is suitable for both sparse and dense video tasks, we compare the efficiency benefits mainly against the more general full self-attention ViViT-L variant, which is utilised by influential follow-up works (e.g., VideoMAE Tong et al. (2022)) over the specialised Factorised Encoder (FE) version, which models temporally only one token per frame, being suitable for classification tasks, but not for dense tasks.
We consider a video with $T$ frames, each frame with $N$ spatial tokens of dimension $D$, and a model with $L$ layers. The full self-attention ViViT flattens the video sequence and computes attention scores between all $T\cdot N$ pairs of tokens followed by a token-wise MLP block, resulting in a complexity $\mathcal{O}(LT^{2}N^{2}D+LTND^{2})$, with the first term associated to self-attention dominating the overall computation. For our hybrid architecture, we consider the complexity of the ViT blocks and the gated LRU blocks separately. The $L$ ViT blocks perform spatial attention over the $N$ tokens within each frame, resulting in $\mathcal{O}(LT(N^{2}D+ND^{2}))$ complexity. The $L$ gated LRU blocks perform temporal processing across the $T$ frames for all $N$ tokens in parallel, contributing $\mathcal{O}(LTND^{2})$ complexity. Hence the total complexity for the proposed model is $\mathcal{O}(LTN^{2}D+LTND^{2})$. The main difference is that ViViT full self-attention scales quadratically with the number of frames ($T^{2}$) and the number of spatial tokens ($N^{2}$), whereas our model scales linearly with the number of frames ($T$) and quadratically with the number of spatial tokens ($N^{2}$).
[…]
@@ 5 Experiments
We present results for supervised video classification and self-supervised masked auto-encoding with frozen representations evaluated on two downstream tasks: video classification and point tracking. To analyse the memory capabilities of our model, we also include a reconstruction task of frames seen in the distant past. Using the same task, we study the generalisation capabilities to longer sequences than seen during training. We follow the ViT scaling configurations and, unless otherwise stated, we use the Base version for our model for all our experiments. We specify the number of parameters for all models considered in our experiments, and we include in the supplementary material all the model configurations, training hyperparameters, and data augmentations used in all experiments.
@@
Datasets: We use mainly two large-scale real-world datasets for ablations and for SOTA comparison on the supervised video classification task. Kinetics400 (Carreira and Zisserman, 2017a) contains 241,512 videos [^2] Kinetics is a dynamic dataset (videos may be removed from YouTube). Our current version has 241,512 videos, compared to 267,000 videos reported in (Arnab et al., 2021), so a decrease of almost 10%, noticeable in the final performance. across train, validation, and test splits, 10s-long (25fps), spanning 400 classes. This dataset is known to require modelling appearance for successful action recognition. To challenge our model’s capability of understanding motion, we also use SSv2 dataset (Goyal et al., 2017), which contains 220,847 shorter videos (2-6s long), sampled at 12fps, representing 174 classes. This dataset includes actions that differ in finer motion-related details, requiring a deeper temporal understanding, e.g. pouring something into something vs pretending to pour something into something.
5.1 Ablations
Gated LRU block: We ran multiple ablations for the components of the gated LRU block used in TRecViT using SSv2 supervised classification as task, see Table 1. We study the impact of the different components and of the initialisation range for the eigenvalues of the recurrent matrix, as mentioned in section 3.1. We report top-1 classification accuracy, i.e. the percentage of videos for which the model’s highest-confidence prediction matches the ground truth label. The results show that the skip connection is essential to successful training; without it, we need to lower the learning rate for the training to take off at all, and the final performance is still very poor. Removing the input and recurrent gates leads to slightly reduced performance, in line with the observations on SSM in language. The input gate adds selectivity on the input, while the recurrent gate modulates how far in the past the model looks. Removing the 1D convolutional layer leads to a bigger decrease, which again is in line with the original Griffin paper that mentioned the important role of the Conv1D layer in extracting local temporal features.
Ablation Top-1(%)
Gated LRU block [0.6, 0.999] (ours) 66.7
Griffin range [0.9, 0.999] 66.2
No skip connection (lower lr) 15.5
No input gate, no recurrent gate 66.0
No convolutional layer 64.9
Table 1: Ablation on Gated LRU components and eigenvalue initialisation. Task: SSv2 classification. Metric: top-1 accuracy (higher is better).
@@ 5.1 Ablations
Training from scratch: We run an experiment to compare the proposed factorisation against simpler adaptations of LRU to video, similar to the ones used in ViViT and VideoMamba, i.e. patchify and flatten the entire video, add spatio-temporal positional embeddings, and apply 1D self-attention or SSM blocks as done in language tasks. Figure 4 compares the proposed TRecViT with these two baselines, with all models being trained from scratch on supervised classification on SSv2. We consider the Small version for all models as the larger Base version shows stability issues when trained from scratch, as reported in other works as well (Li et al., 2024; Arnab et al., 2021). As expected, the performance on this challenging dataset when training from scratch is far from SOTA, but it clearly shows that the proposed factorisation has superior video modelling capabilities compared to baselines, ViViT-S with full self-attention being the closest competitor. PureLRU’s performance is very poor, which is in line with the findings from VideoMamba that bidirectional (non-causal) processing of the input is needed for good performance.
Same factorisation, different temporal blocks: We compare our proposed hybrid architecture against other factorised variations that have similar efficiency; see Figure 5 in the appendix for memory footprint and FLOPs comparison. The results are included in Table 2 and clearly show that the proposed hybrid architecture is superior in performance and in training throughput compared to architectures using LSTM or factorised self-attention to integrate temporal information. The model using Conv1D as temporal module is faster to train, but its accuracy is far from competitive.
Additional ablations: We include in the appendix ablations by running supervised classification on SSv2 using different hyperparameter values: temporal dimension of video patches, minimal radius for eigenvalue initialisation, window size for the 1D convolution in LRU. We also include an experiment considering multiple seeds and report mean and variance.
Temporal module (w skip) Top-1 (%) Steps per sec (training)
ViViT factorised self-attention 62.4 8.3
ViT-LSTM 63.7 6.0
ViT-Conv1D 40.3 10.5
TRecViT 66.7 8.9
Table 2: Comparison of different temporal modules used in hybrid setups similar to our proposed TRecViT. Task: SSv2 classification. Metric: Top-1 accuracy (higher is better).
5.2 SOTA comparison on supervised video classification
We compare the performance of TRecViT against top models in the literature, considering both causal and non-causal architectures; see Table 3 for SSv2 results and Table 4 for Kinetics400 results. On SSv2, our model achieves state-of-the-art performance compared to causal baselines (TSM, RViT, causal ViViT) and outperforms or is competitive with all non-causal baselines. Notably, it outperforms the popular ViViT-L by 2.3% despite having about $3\times$ less parameters.
On Kinetics400, TRecViT outperforms convolutional architectures (I3D) and some Transformer-based architectures (TimeSformer, ViViT-B, causal ViViT-L). Compared to RViT and ViViT-L, the performance is competitive, but slightly lower. We attribute this in part to the reduced number of videos in the dataset as mentioned above – when training both TRecViT and ViViT-L on the same number of videos, the performance gap is significantly reduced. In addition, this result could also reflect the difference between the two datasets mentioned above (appearance vs motion) and highlights that TRecViT is superior at modelling motion compared to baselines, but on Kinetics400 where the appearance is enough for successful classification, the performance is on par.